What is the probability that a market inspector will discover at least three violations of the public health code

1
Expert's answer
2022-02-25T11:07:29-0500

The probability that a market inspector will discover at least three violations of the public health code is given by, "P(X \ge 3) =1-p(X\lt3)=1-\{ P(X=0)+P(X=1)+P(X=2)\}"